OpenStudy (naga447):

could someone help me find the slope of this equation? I will give a medal for who ever helps! g(x) = 8x - 25

OpenStudy (solomonzelman):

y = mx + b where m is the slope

OpenStudy (bradely):

g(x) = 8x - 25 it is in the form y =mx+c slope 8 source: http://www.mathskey.com/question2answer/

OpenStudy (naga447):

Okay thank you both!